x^3dx+(y+1)^2dy=0


1
Expert's answer
2022-03-01T18:43:04-0500

Let us solve the differential equation

"x^3dx+(y+1)^2dy=0."

It follows that

"\int x^3dx+\int(y+1)^2dy=C," and we conclude that the general solution of this differential equation is

"\frac{x^4}4+\frac{(y+1)^3}3=C."